OpenStudy (anonymous):

the fundamental property of rational expressions 2/5y

Directrix (directrix):

Thanks for this post. I had not heard the Fundamental Property of Rational Expressions by this name. ---------------- It says: Fundamental property of rational expressions- If K represents a factor then (PK)/(QK) = P/Q where K not equal to zero. This property used in simplifying the rational expression to the lowest terms. Example: (3x - 6)/( 4x-8) = {3(x - 2)}/{4(x - 2)} = 3/4 Question: Is your problem (2/5)* y or is it 2/(5y)? Thanks.